determine the oxidation states of the elements in the compounds listed. none of the oxygen-containing compounds are peroxide’s or superoxides.
a.H2SO4
b.Mg(OH)2
c.P4O6
1
Expert's answer
2019-10-07T06:41:03-0400

For "H_2So_4" :

"Oxidation" state of Hydrogen = "+1"

Oxidation state of oxygen = "-2"

And,

"({(1×2)+(-2×4)})+O.S.\ of \ S=0\\so,\\ oxidation\ state\ of\ sulphur=+6"


For "Mg(OH)_2:"

Oxidation state of Oxygen "=-2"

Oxidation state of hydrogen "=+1"

And,

"O.S.\ of\ Mg+(2×-1)=0\\"

So,

Oxidation state of magnesium "=+2"


For "P_4O_6:"

Oxidation state of Oxygen "=-2"

And ,

"4\times O.S.\ of\ P+(-2×6)=0"

So,

Oxidation state of phosphorus "=+3"